OpenStudy (nsh4267):

A mixture of 50 liters of paint is 25% red tint, 30% yellow tint and 45% water. 6 liters of yellow tint are added to the original mixture. The percent of yellow tint in the new mixture is

OpenStudy (sparrow2):

it will be the same 30%

OpenStudy (nsh4267):

thats not right

OpenStudy (study312):

A mixture of 40 liters of paint is 25% red tint, 30% yellow tint and 45% water. 10 liters of yellow tint are added to the original mixture. The percent of yellow tint in the new mixture is? Since 30% of 40 liters=12 liters of yellow tint 10 +40=50 Total mixture, but since we had 12 liters of yellow tint + 10 new liters=22 liters of yellow tint 22/50=44% of the new mixture is yellow tint. - http://web2.0calc.com/questions/a-mixture-of-40-liters-of-paint-is-25-red-tint-30-yellow-tint-and-45-water-10-liters-of-yellow-tint-are-added-to-the-original-mixture

OpenStudy (study312):

Just adjust that to 6 liters of yellow paint because that answer is 10

OpenStudy (study312):

0.30 • 50 = ? can u tell me?

OpenStudy (nsh4267):

15

OpenStudy (study312):

6 + 50 = ?

OpenStudy (nsh4267):

56

OpenStudy (study312):

Good!

OpenStudy (study312):

6 + 15 = ?

OpenStudy (nsh4267):

21

OpenStudy (nsh4267):

so 100*22/56

OpenStudy (study312):

Yes! 21 ÷ 56 = ? it's 21 not 22

OpenStudy (study312):

21 ÷ 56

OpenStudy (nsh4267):

thats what i meant which equals 37.5

OpenStudy (study312):

37.5 is your answer

OpenStudy (nsh4267):

thank you!

OpenStudy (study312):

Np

OpenStudy (study312):

was this multiple choice?

OpenStudy (nsh4267):

no